Provide your answers in the space provided or the back of the page. Make sure to scan all work pages. Any quantitative questions require showing your work for full credit.
Round all $ problems to the nearest cent.
All other calculations must be to at least the fifth decimal place.

1) The following table contains the CPI and the Federal Minimum hourly wage rates. (Show work for credit)

YEAR Hourly Min. Wage, CPI

1945          $0.40        18
1955          0.75         26.8
1965           1.25        31.5
1975           2.10        53.8
1985           3.35        107.6
1995            4.25       152.4
2005            5.15       195.3
2015            7.25       233.8
2018                        252.0

a.) Convert each nominal minimum wage into bae year dollars and rank them (by year) from lowest to highest real wage.

b.) Determine how much the current minimum wage must be to equal the purchasing power of the highest real minimum wage on the list.

c.) Determine how much the current minimum wage must be to equal the purchasing power of the lowest real minimum wage on the list.
Answer #1

a) We first normalize the price index by 1945 value and calculate the real wage as = nominal wage/normalized price index.

Year Min wage CPI Normalized CPI Real Wage Rank
1945 0.4 18 1 0.4 1
1955 0.75 26.8 1.49 0.5 3
1965 1.25 31.5 1.75 0.714 8
1975 2.1 53.8 3 0.7 7
1985 3.35 107.6 6 0.56 5
1995 4.25 152.4 8.47 0.5 4
2005 5.15 195.3 10.85 0.475 2
2015 7.25 233.8 13 0.58 6

b) The highest real min wage is 0.714(1965). The current normalized CPI is 252/18 = 14.

To match highest real min wage, Current min wage should be 0.714*14 = $ 10.

c) Also the lowest real min wage is 0.4(1945). The current normalized CPI is 252/18 = 14.

To match lowest real min wage, Current min wage should be 0.4*14 = $ 5.6